# Dynamics of Pseudoentanglement

### v1

@@ Appendix F Upper bound on relaxation time with local gate set
In Sec. IV.3 we study the relaxation time of the Markov chains induced by the local automaton gate set $\mathcal{G}=\{\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$ and, based on numerical observations, we conjecture the scaling $t_{\rm rel}^{(m)}=\Theta(N)$, independent of $m$ (Conjecture 1). While the single-particle spectrum, solved analytically in App. D.1, gives a lower bound $t_{\rm rel}^{(m)}\geq\Omega(N)$, an upper bound is harder to obtain analytically. Here we prove the upper bound
$t_{\rm rel}^{(m)}\leq O(mN^{5}).$ (105)
We believe this is a very loose upper bound (again Conjecture 1 corresponds to $t_{\rm rel}^{(m)}\leq O(N)$) and it would be interesting to identify different techniques to tighten it.
Our proof technique is based on the comparison method Levin and Peres (2017), which provides a bound
$t_{\rm rel}\leq B\tilde{t}_{\rm rel}$ (106)
on the relaxation time $t_{\rm rel}$ of a Markov chain $\Gamma$ in terms of the relaxation time $\tilde{t}_{\rm rel}$ of another Markov chain $\tilde{\Gamma}$ on the same space. To define the coefficient $B$, we first need to introduce some background and notation. We consider a graph $(V,E)$ with vertices $V$ and edges $E$; the reversible Markov chain $\Gamma$ allows transitions between any two vertices connected by an edge, with probability $\Gamma_{e_{1}e_{2}}$ where $(e_{1},e_{2})=e\in E$ denotes the edge. We also have another graph with the same vertices but different edges, $(V,\tilde{E})$, with a Markov chain $\tilde{\Gamma}$ allowing transitions along the $\tilde{E}$ edges. We assume that, for any edge $\tilde{e}=(x,y)\in\tilde{E}$ in the second graph, one can find a path $\gamma_{xy}$ along the first graph $(V,E)$ connecting $x$ to $y$. $|\gamma_{xy}|$ denotes the length of that path. We say that $e\in\gamma_{xy}$ if the path traverses edge $e\in E$. With this notation in place, we can write down the expression for the coefficient $B$ in the comparison bound:
@@
Our strategy is to use our Markov chain (induced by the local automaton gate set $\mathcal{G}$) as $\Gamma$ and use the Markov chain induced by ‘simple permutations’ Gowers (1996); Hoory et al. (2005) as $\tilde{\Gamma}$. ‘Simple permutations’ are defined as the gate set $\tilde{\mathcal{G}}=\{v_{ijk,\sigma}:\,i,j,k\in[N],\,\sigma\in S_{8}\}$, where $v_{ijk,\sigma}$ acts as a 3-qubit automaton gate [^14] In ‘width-2’ simple permutations Hoory et al. (2005); Brodsky and Hoory (2008), which we use here to establish our upper bound, $\sigma$ is restricted to act as a flip of bit $i$ controlled by an arbitrary Boolean function of bits $j,k$. The Toffoli gate falls in this category, with the Boolean function being AND. (associated to a permutation $\sigma\in S_{8}$) on qubits $i,j,k$. Note that the choice of qubits $i,j,k$ is also arbitrary, in particular they need not be consecutive. So ‘simple permutations’ are 3-local but not geometrically local. ‘Simple permutations’ have been studied in the probability literature Gowers (1996); Hoory et al. (2005) due to their good mixing properties, and several bounds on their relaxation and mixing times are known. For our purpose, we will need the following fact:
@@
‘Simple permutations’ $\tilde{\mathcal{G}}$ induce a Markov chain on $\Sigma_{m}$ whose relaxation time is bounded above as
@@
Ref. Brodsky and Hoory (2008) considers the random walk induced by ‘simple permutations’ on ordered $m$-tuples, $\mathbf{z}=(z_{1},\dots z_{m})$, rather than sets $S=\{z_{1},\dots z_{m}\}\in\Sigma_{m}$. It shows that the former has relaxation time $\tilde{t}_{\rm rel}^{(m,\text{ord.})}\leq O(mN^{2})$. We will show that
@@
Our choice of paths $\gamma_{S,S^{\prime}}$ (see proof of Eq. (113) above) depends only on the gate $v$ (or more precisely, on one of the $\mu_{S,S^{\prime}}$ possible gates that map $S$ to $v(S)$, chosen arbitrarily): we decompose $v=u_{l}u_{l-1}\cdots u_{1}$ into a sequence of $l\leq O(N)$ local automaton gates, $u_{i}\in\mathcal{G}$. The condition $\gamma_{S,v(S)}\ni e$ thus requires that $e_{1}=u_{i}\cdots u_{1}(S)$ and $e_{2}=u_{i+1}(e_{1})$ for some $i\in[l]$; equivalently, it requires $S=u_{1}\cdots u_{i}(e_{1})$. Therefore, given $e$ and $v$, there are $O(N)$ choices for $S$, indexed by the integer $i$. The $|\tilde{\mathcal{G}}|$ choices for $v$ cancel out the denominator in Eq. (116), and we obtain the desired bound. ∎
Putting it all together, Fact F.1 bounds the relaxation time in the ‘simple permutation’ Markov chain as $\tilde{t}_{\rm rel}^{(m)}\leq O(mN^{2})$; Facts F.2, F.3, and F.4 together yield the comparison coefficient $B\leq O(N^{3})$; plugging these into the comparison bound (Eq. (106)) gives the desired result (Eq. (108)).
We close with two observations. First, since our best lower bound on the mixing time (for general $m$) is $t_{\rm mix}^{(m)}\geq\tilde{\Omega}(mN)$, this upper bound on $t_{\rm rel}^{(m)}$ is not sufficient to separate the relaxation and mixing times. Second, using the relaxation time, we can also bound the mixing time above by Eq. 28. This gives the (likely loose) upper bound $t_{\rm mix}^{(m)}\leq O(m^{2}N^{6})$.
Appendix G Upper bound on pseudothermalization time for subset states
In this Appendix we prove the upper bound Eq. (44) on the pseudothermalization time for subset states, Sec. V. We start from the key result, Eq. (38-39), which relates pseudothermalization to the equilibration of the Markov chains $\Phi_{m+\delta\leftarrow K}[p_{t}]$ on $\Sigma_{m+\delta}$, for $\delta=0,\dots m$.
@@ Appendix G Upper bound on pseudothermalization time for subset states
$O_{\delta}$ $=\binom{K}{m}^{-1}\binom{K}{m+\delta}\times$
×$\quad\times\sum_{\begin{subarray}{c}S^{\prime},S^{\prime\prime}\in\Sigma_{m}:\\ |S^{\prime}\setminus S^{\prime\prime}|=\delta\end{subarray}}\ket{S^{\prime}}f^{(m+\delta)}_{p_{t}}(S^{\prime}\cup S^{\prime\prime})\bra{S^{\prime\prime}},$ (120)
where $f^{(m+\delta)}_{p_{t}}$ is defined in Eq. (40). The diagonal term $O_{0}$ provides a lower bound: $\|\mathcal{M}^{(m)}_{p_{t}}\|\geq\|O_{0}\|$, which is used to derive the lower bound on the prethermalization time, Eq. (43). To get an upper bound, we need to consider the off-diagonal terms as well.
@@
$\|A\|_{\rm tr}=\|\sum_{ij}A_{ij}\outerproduct{i}{j}\|_{\rm tr}\leq\sum_{ij}\|A_{ij}\outerproduct{i}{j}\|_{\rm tr}=\sum_{ij}|A_{ij}|.$ | . (122)
@@
$=\binom{m}{\delta}\binom{K-m}{\delta}\Delta_{m+\delta},$ (123)
@@
$\Delta_{m+\delta}=\|\Phi_{m+\delta\leftarrow K}[p_{t}]-\pi_{m+\delta}\|_{\rm tv}.$ (124)
@@
$\|\mathcal{M}^{(m)}_{p_{t}}\|_{\rm tr}$ $\leq\Delta_{2m}\sum_{\delta=0}^{m}\binom{m}{\delta}\binom{K-m}{\delta}=\binom{K}{m}\Delta_{2m}.$ (125)
Thus, to make the LHS smaller than $\epsilon$, the Markov chain on $\Sigma_{2m}$ must equilibrate to within total variation error $\Delta_{2m}\leq\binom{K}{m}^{-1}\epsilon$. Defining the mixing time $t_{\rm mix}(\epsilon)$ based on a total variation threshold $\epsilon$ (see Sec. II.2), one has the following general inequality:
$t_{{\rm mix}}(2^{-l-1})\leq lt_{{\rm mix}}(1/4).$ (126)
It follows then that
$t_{\rm PTherm}^{(m)}\leq\left[\log_{2}\binom{K}{m}\right]t_{\rm mix}^{(2m)}\leq m\log_{2}(K)t_{\rm mix}^{(2m)}.$ (127)
@@
$t_{\rm PTherm}^{(m)}\leq O(\log(K)m^{2}N^{2}).$ (128)
In the case of $K=\exp({\rm polylog}(N))$ (where the entropy of the random subset ensemble is polylogarithmic) this reduces to
$t_{\rm PTherm}^{(m)}\leq\tilde{O}(m^{2}N^{2}),$ (129)
where the symbol $\tilde{O}$ represents an upper bound up to $\log(N)$ factors.

### v3

@@ Appendix F Upper bound on relaxation time with local gate set
In Sec. 4.3 we study the relaxation time of the Markov chains induced by the local automaton gate set $\mathcal{G}=\{\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$ and, based on numerical observations, we conjecture the scaling $t_{\rm rel}^{(m)}=\Theta(N)$, independent of $m$ (Conjecture 1). While the single-particle spectrum, solved analytically in App. D.1, gives a lower bound $t_{\rm rel}^{(m)}\geq\Omega(N)$, an upper bound is harder to obtain analytically. Here we prove the upper bound
$t_{\rm rel}^{(m)}\leq O(mN^{5})$ (92)
based on the comparison method levin_markov_2017 applied to “simple permutations” gowers_almost_1996 ; brodsky_simple_2008 , a set of 3-local but not geometrically local reversible classical gates202020 Based on the results of Ref. he_pseudorandom_2024 , which appeared after the first version of this work, the comparison method applied to a set of geometrically-local 3-bit classical gates known as DES[2] (which is a super-set of our gate set $\mathcal{G}$) would yield an improved bound $t_{\rm rel}\leq O(mN)$. . This is very loose compared to Conjecture 1 ($t_{\rm rel}\leq O(N)$) but serves to rigorously bound above [via Eq. (29)] the mixing time of our Markov chains, and thus the $m$-copy indistinguishability times $t_{\rm indist}^{(m)}$, as $O(m\,{\rm poly}(N))$.
The comparison method bounds the relaxation time $t_{\rm rel}$ of a Markov chain $\Gamma$ in terms of the relaxation time $\tilde{t}_{\rm rel}$ of another Markov chain $\tilde{\Gamma}$ on the same space as levin_markov_2017
$t_{\rm rel}\leq B\tilde{t}_{\rm rel},$ (93)
in terms of a “congestion ratio” $B$. To define $B$, we first need to introduce some background and notation. We consider a graph $(V,E)$ with vertices $V$ and edges $E$; the reversible Markov chain $\Gamma$ allows transitions between any two vertices connected by an edge, with probability $\Gamma_{e_{1}e_{2}}$ where $(e_{1},e_{2})=e\in E$ denotes the edge. We also have another graph with the same vertices but different edges, $(V,\tilde{E})$, with a Markov chain $\tilde{\Gamma}$ allowing transitions along the $\tilde{E}$ edges. We assume that, for any edge $\tilde{e}=(x,y)\in\tilde{E}$ in the second graph, one can find a path $\gamma_{xy}$ along the first graph $(V,E)$ connecting $x$ to $y$. $|\gamma_{xy}|$ denotes the length of that path. We say that $e\in\gamma_{xy}$ if the path traverses edge $e\in E$. With this notation in place, we can write down the expression for the congestion ratio $B$:
@@
Our strategy is to use our Markov chain (induced by the local automaton gate set $\mathcal{G}$) as $\Gamma$ and use the Markov chain induced by ‘simple permutations’ gowers_almost_1996 ; hoory_simple_2005 as $\tilde{\Gamma}$. ‘Simple permutations’ are defined as the gate set $\tilde{\mathcal{G}}=\{v_{ijk,\sigma}:\,i,j,k\in[N],\,\sigma\in S_{8}\}$, where $v_{ijk,\sigma}$ acts trivially on all bits not in $\{i,j,k\}$, and as a permutation $\sigma\in S_{8}$ on bits $\{i,j,k\}$. We further restrict $\sigma\in S_{8}$ to be a ‘width-2’ simple permutation, i.e. a permutation that leaves two of the three input bits unchanged hoory_simple_2005 ; brodsky_simple_2008 (the Toffoli gate falls in this category, so $\mathcal{G}\subset\tilde{\mathcal{G}}$). Note that the bits $i,j,k$ need not be consecutive, so ‘simple permutations’ are 3-local but not geometrically local. ‘Simple permutations’ have been studied in the probability literature gowers_almost_1996 ; hoory_simple_2005 due to their good mixing properties, and several bounds on their relaxation and mixing times are known. For our purpose, we will need the following fact:
@@
‘Width-2 simple permutations’ $\tilde{\mathcal{G}}$ induce a Markov chain on $\Sigma_{m}$ whose relaxation time is bounded above as
@@
Ref. brodsky_simple_2008 considers the random walk induced by ‘simple permutations’ on ordered $m$-tuples, $\mathbf{z}=(z_{1},\dots z_{m})$, rather than sets $S=\{z_{1},\dots z_{m}\}\in\Sigma_{m}$. It shows that the former has relaxation time $\tilde{t}_{\rm rel}^{(m,\text{ord.})}\leq O(mN^{2})$. We will show that
@@
Our choice of paths $\gamma_{S,S^{\prime}}$ (see proof of Eq. (100) above) depends only on the gate $v$: we decompose $v=u_{l}u_{l-1}\cdots u_{1}$ into a sequence of $l\leq O(N)$ local automaton gates, $u_{i}\in\mathcal{G}$. The condition $\gamma_{S,v(S)}\ni e$ thus requires that $e_{1}=u_{i}\cdots u_{1}(S)$ and $e_{2}=u_{i+1}(e_{1})$ for some $i\in[l]$; equivalently, it requires $S=u_{1}\cdots u_{i}(e_{1})$. Therefore, given $e$ and $v$, there are $O(N)$ choices for $S$, indexed by the integer $i$. The $|\tilde{\mathcal{G}}|$ choices for $v$ cancel out the denominator in Eq. (103), and we obtain the desired bound. ∎
Putting it all together, Fact F.1 bounds the relaxation time in the ‘simple permutation’ Markov chain as $\tilde{t}_{\rm rel}^{(m)}\leq O(mN^{2})$; Facts F.2, F.3, and F.4 together yield the congestion ratio $B\leq O(N^{3})$; plugging these into the comparison bound [Eq. (93)] gives the desired result [Eq. (95)].
Appendix G Upper bound on $m$-copy indistinguishability time for subset states
In this Appendix we prove the upper bound Eq. (45) on the $m$-copy indistinguishability time for subset states, Sec. 5. We start from the key result, Eq. (39-40), which relates ensemble pseudothermalization to the equilibration of the Markov chains $\Phi_{m+\delta\leftarrow K}[p_{t}]$ on $\Sigma_{m+\delta}$, for $\delta=0,\dots m$.
@@ Appendix G Upper bound on $m$-copy indistinguishability time for subset states
$O_{\delta}$ $=\binom{K}{m}^{-1}\binom{K}{m+\delta}\sum_{\begin{subarray}{c}S^{\prime},S^{\prime\prime}\in\Sigma_{m}:\\ |S^{\prime}\setminus S^{\prime\prime}|=\delta\end{subarray}}\ket{S^{\prime}}f^{(m+\delta)}_{p_{t}}(S^{\prime}\cup S^{\prime\prime})\bra{S^{\prime\prime}},$ (105)
where $f^{(m+\delta)}_{p_{t}}$ is defined in Eq. (41). The diagonal term $O_{0}$ provides a lower bound: $\|\mathcal{M}^{(m)}_{p_{t}}\|_{\rm tr}\geq\|O_{0}\|_{\rm tr}$, which is used to derive the lower bound on the $m$-copy indistinguishability time, Eq. (44). To get an upper bound, we need to consider the off-diagonal terms as well.
@@
$\|A\|_{\rm tr}=\left\|\sum_{ij}A_{ij}\outerproduct{i}{j}\right\|_{\rm tr}\leq\sum_{ij}\|A_{ij}\outerproduct{i}{j}\|_{\rm tr}=\sum_{ij}|A_{ij}|.$ | . (107)
@@
$=2\binom{m}{\delta}\binom{K-m}{\delta}\Delta_{m+\delta},$ (108)
@@
$\Delta_{m+\delta}=\frac{1}{2}\sum_{S\in\Sigma_{m+\delta}}|\Phi_{m+\delta\leftarrow K}[p_{t}](S)-\pi_{m+\delta}(S)|.$ (109)
@@
$\frac{1}{2}\|\mathcal{M}^{(m)}_{p_{t}}\|_{\rm tr}$ $\leq\Delta_{2m}\sum_{\delta=0}^{m}\binom{m}{\delta}\binom{K-m}{\delta}=\binom{K}{m}\Delta_{2m}.$ (110)
Thus, to make the left hand side smaller than $\epsilon$, the Markov chain on $\Sigma_{2m}$ must equilibrate to within total variation error $\Delta_{2m}\leq\binom{K}{m}^{-1}\epsilon$. Defining the mixing time $t_{\rm mix}(\epsilon)$ based on a total variation threshold $\epsilon$ (see Sec. 2.2), one has the following general inequality:
$t_{{\rm mix}}(\epsilon=2^{-l-1})\leq lt_{{\rm mix}}(\epsilon=1/4).$ (111)
It follows then that (taking the mixing time based on $\epsilon=1/4$ as usual)
$t_{\rm indist}^{(m)}\leq\left[\log_{2}\binom{K}{m}\right]t_{\rm mix}^{(2m)}\leq m\log_{2}(K)t_{\rm mix}^{(2m)}.$ (112)
@@
$t_{\rm indist}^{(m)}\leq O(\log(K)m^{2}N^{2}).$ (113)
In the case of $\log K={\rm polylog}(N)$ (where the entropy of the random subset ensemble is polylogarithmic) this reduces to
$t_{\rm indist}^{(m)}\leq\tilde{O}(m^{2}N^{2}),$ (114)
where the symbol $\tilde{O}$ represents an upper bound up to ${\rm polylog}(N)$ factors.